The base of an isosceles triangle is 4 in. less than the length of the equal sides. If the perimeter of the triangle is 38 in., find the length of each of the sides.
HELP!!! :)

Answer by Earlsdon(6103)

Find the lengths of the sides of an isosceles triangle.
Let the length of the equal sides be x inches each.
Then the length of the base is x-4 inches.
The perimeter, P = x + x + (x-4)
x + x + (x-4) = 38 Collect like-terms:
3x-4 = 38 Add 4 to both sides.
3x = 42 Divide both sides by 3.
x = 14 inches for each of the equal sides.
The base is 14-4 = 10 inches.